40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5198     Entry time: Thu Aug 11 18:30:40 2011
Author: Koji 
Type: Update 
Category: SUS 
Subject: ITMY OSEM adjustment 

[Jamie, Koji]

ITMX OSEMs were adjusted so as to have the right DC numbers and the more uniform response to POS excitation.

It is waiting for the free-swinging test.

- ITMX was moved from its position to the north side of the table.

- The table was rebalanced.

- We found that the output of the LR OSEM has an excess noise compared with the other OSEMs.
We tried to swap the LR and SD OSEMs, but the SD OSEM (placed at the LR magnet) showed
the same excess noise at around 10-50Hz.

- We found that one of the EQ stops was touching the mirror. By removing this friction, all of the OSEMs
come to show similar power spectra. Good!

 - Then we started to use LOCKIN technique to measure the sensitivity of the OSEMs to the POS excitation.

Originally the response of the OSEMs was as follows
UL 3.4 UR 4.3 
LL 0    LR 2.5   

After the adjustment of the DC values, final values became as follows
UL 3.9 UR 4.4
LL 3.9  LR 3.2

- We decided to close the light door.

ELOG V3.1.3-